I'm trying to estimate the mean of a population, so I want to take a sample which (regardless of sample size) will have an SD s of 18.4, let's say I want an 86% confidence interval, but I also want my margin of error to be less than 6, how large of a sample do I need to take? ______
Answer #1

Solution :

Given that,

standard deviation =s =   =18.4

Margin of error = E = 6

At 86% confidence level

Z/2 =1.476

sample size = n = [Z/2* / E] 2

n = ( 1.476 * 18.4 / 6 )2

n =20.488

Sample size = n =20

if we take z value 2 decimal so answer is

n = ( 1.48 * 18.4 / 6 )2

n =21